# Optimizing Multi-Modal Trackers via Sensitivity-aware Regularized Tuning

### v1

@@ III-B Modeling Sensitivity for Multi-Modal Tracking
Manifold hypothesis. As shown in Fig. 3, the parameter set $\{\theta\}_{L=L_{c}}$, with certain loss $L_{c}$, does not fill the entire space $\mathbb{R}^{|\theta|}$. Instead, it usually lies in a manifold $\mathcal{M}_{L_{c}}$, which is also continuous in $\mathbb{R}^{|\theta|}$.
Based on the aforementioned manifold hypothesis, to avoid the losing of generalization ability for fine-tuned models on downstream domains, the critical thing lies on optimize the model on local tangent space, shown as the red arrow of the Fig. 3. Thus, in the following section, we investigate the method to approximate and optimize the model on such tangent space.
@@ III-B1 Safeguarding Generalization via Prior Sensitivity
This generalization gap is measured as the weight distance within the Riemannian manifold defined by the FIM. Through geometric perspective, the FIM yields insights that large deviations from pre-trained tangent space entail a higher risk of generalization degradation [45, 46]. Therefore, the FIM serves as a natural and principled prior sensitivity to reflect the erosion of pre-trained knowledge.
Figure 4: The operation-wise eigen-decomposition FIM map (i.e., prior sensitivity) of the pre-trained OSTrack. The high prior sensitivity indicates a tendency to deviate from the pre-trained tangent space, reflecting the disruption of pretrained knowledge.
Eigen-decomposition FIM Approximation. Directly computing the FIM for a large model is computationally intractable due to its $\mathcal{O}\left(|\theta|^{2}\right)$ complexity. Instead, we propose constructing an eigen-decomposition approximation of the FIM without explicitly forming it in full. Focusing on leading eigenvalues captures key tangent geometrics of affecting generalization while reducing complexity. Specifically, we partition the model’s parameters into $N$ disjoint operation groups, $\theta=\left\{\theta^{1},\cdots,\theta^{N}\right\}$ (e.g. MLPs of FFNs and attention QKVs). Assuming inter-operation independence, we yield a group-diagonal FIM: $\mathcal{F}^{(\theta_{0})}=\mathrm{diag}(\mathcal{F}^{(\theta_{0}^{1})},\dots,\mathcal{F}^{(\theta_{0}^{N})})$, where $\mathcal{F}^{(\theta_{0}^{j})}\in\mathbb{R}^{|\theta^{j}|\times|\theta^{j}|}$ corresponds an independent parameter group $j$. Parameters from different groups thus span orthogonal subspaces with no second-order coupling.
Empirical studies [47, 48] have shown that the FIM of deep networks are spectrally concentrated, featuring a few large eigenvalues and a long, flat tail, which motivates a low-rank approximation using its leading eigenpairs. Let $\tilde{\mathcal{F}}^{(\theta_{0}^{j})}={(V^{j})}\Lambda^{j}{(V^{j}})^{T}$ be the eigen-decomposition, with unit eigenvectors $V^{j}=[v^{j}_{1},\dots,v^{j}_{K}]$ and eigenvalues $\Lambda^{j}=\mathrm{diag}(\lambda^{j}_{1},\dots,\lambda^{j}_{K})$ sorted $\lambda^{j}_{1}\geq\cdots\geq\lambda^{j}_{K}\geq 0$.
@@
See Appendix.
Prior Sensitivity Measure. Specifically, we estimate large eigenvalues of FIM using symmetric finite-difference probes [50]. Around $\theta_{0}$ on $D_{0}$, small fixed-radius perturbations expose the directions that maximize the normalized generalization gap, which is equivalent to Rayleigh quotient of FIM. This procedure systematically maps the pre-trained loss landscape to recover the sharpest generalization shifts and identify the most sensitive parameters. For each group $\theta^{j}$, by sampling various unit directions $\epsilon^{j}\in\mathbb{R}^{|\theta^{j}|}$ from an isotropic Gaussian distribution $\epsilon^{j}\sim\mathcal{N}\left(\mathbf{0},I\right)$ on the weight sphere, we obtain the empirical measures:
@@ III-B2 Stabilizing Adaptation via Transfer Sensitivity
Figure 5: The instantaneous transfer sensitivity on the VisEvent dataset observed during fine-tuning. The high sparsity in the sensitivity map indicates that only a few gradients (light areas) dominate the updates.
@@
where $\delta$ is a random perturbation centered on the current update step, i.e., $\delta\sim\mathcal{N}\left(\alpha\mathcal{G},\sigma^{2}I\right)$, $\alpha$ is the learning rate and $\sigma$ controls the noise. This denotes that updates follow the gradient with inherent noise (due to stochastic sampling, batch gap, etc.).
@@
$\varepsilon_{ada}\approx\alpha\sqrt{2\sigma^{2}}\left\|\mathcal{G}\right\|_{2}.$ (12)
Thus, we can deduce that sparse gradients exacerbate tuning instability and adaptation risk. This highlights the need to avoid concentrating updates on specific parameters and emphasizes the importance of a balanced transfer gradient.
@@
$s_{n}^{t}=(\frac{\partial\mathcal{L}\left(\theta\mid M\right)}{\partial\theta_{n}})^{2},$ (13)
@@
1Input: pre-trained dataset $D_{0}$, downstream task dataset $D_{t}$, pre-trained weight $\theta_{0}$; number of eigenvalues $K$, number of model operations NN, training steps $T$; initialized eigenvalue set $\Lambda$;
2Output: Optimal multi-modal tracker parameters $\theta_{t}$;
3Step1: Prior Sensitivity Estimation:
@@
4 Calculate prior sensitivity $s_{j}^{p}$ via Eq. (8);
5Step2: Transfer Sensitivity Estimation and Regularized Tuning:
@@
6  Compute loss $\mathcal{L}(\theta\mid M_{i})$ and gradients;
7  Update transfr sensitivity by $s_{n}^{t}=(\frac{\partial\mathcal{L}\left(\theta\mid M_{i}\right)}{\partial\theta_{n}})^{2}$;
8  Normalize and combine sensitivity $s_{n}$ by Eq. (14);
9  Parameter update via Eq. (15);
@@ 6  Compute loss $\mathcal{L}(\theta\mid M_{i})$ and gradients;
Algorithm 1 Sensitivity-Aware Regularization Tuning
III-C Sensitivity-Aware Regularization Tuning
We regularize the learning process by previously discussed sensitivity metrics to derive enhanced trackers. To align with the unified regularization tuning, we rank both the operation-wise prior sensitivity and parameter-wise transfer sensitivity, and then normalize them within the continuous range of $[0,1]$. To harmonize these two sensitivities, we devise a dynamic linear schedule to adjust their weighted combination. At the beginning of training, prior sensitivity plays a dominant role, contributing a weight of $\kappa$ (where $\kappa\in[0,1]$). As training progresses, the influence of transfer sensitivity gradually increases, reaching the same weight $\kappa$ by the end. This ensures that the model initially focuses on retaining pre-trained knowledge and progressively shifts to emphasize training stability. Formally, at each training step $t$, the combined sensitivity is updated as follows:
$s_{n}=(\kappa+(1-2\kappa)\frac{t}{T})s_{j}^{p}+(1-\kappa-(1-2\kappa)\frac{t}{T})s_{n}^{t},$ (14)
@@ III-C Sensitivity-Aware Regularization Tuning
where $\alpha$ is the learning rate. This formulation suggests that more sensitive parameters should retain their previous states to a greater extent, to avoid oscillations or over-adjustments. The detailed regularization tuning process is given in Algorithm.1. We set the sensitivity harmony coefficient $\kappa=0.6$ in our experiments. Further comprehensive analysis can be found in Tab. VI.
@@ III-D Learning Objectives
Figure 6: Visualization of the precision and success plots of the FE108, VisEvent, CoeSot and LasHeR datasets. Zoom in to see details.
@@ IV-A2 Evaluation Metrics
Adhering to recognized standards [16, 17, 1, 11], we evaluated tracking performance with following metrics. For all RGB-Event benchmarks (FE108, VisEvent and CoeSot), We utilize two widely used metrics, i.e., success rate (SR), precision rate (PR). For DepthTrack, precision (Pr) and recall (Re) are used, with F-score ($F=\frac{2Re\cdot Pr}{Re+Pr}$) as the primary measure. For VOT-RGBD2022, we assess accuracy (Acc), robustness (Rob), and expected average overlap (EAO). For LasHeR, success rate (SR) and precision rate (PR) are employed. For RGBT234, maximum success rate (MSR) and maximum precision rate (MPR) served as evaluation metrics. Notably, our SR/PR precisely aligns with SR/PR in ViPT and OneTrack, Suc/Pre in SDSTrack, Success/Precision in UnTrack, or AUC/P in SUTrack. For all metrics, the larger, the better.
@@ IV-A3 Pre-trained Models and Compared Methods
In this paper, we choose three prototypical one-stream RGB-based trackers as pre-trained baselines: OSTrack [5], the most widely adopted model; DropTrack [40] and SUTrack [8], which offer improved generalization. Notably, OSTrack and DropTrack are built upon ViT-B/16 [57], while SUTrack adopts the HiViT-B/16 [58] as the backbone. Following their pre-trained configurations, these models differ in input resolution: OSTrack ($\times128$$\times256$), DropTrack and SUTrack ($\times192$$\times384$) for template and search regions, respectively. To comprehensively validate the effectiveness of our method, we conduct the following experiments in Tab. I and Tab. II. First, we construct strong RGB-based single-modal baselines under full fine-tuning. Then, we evaluate our method under both full and parameter-efficient fine-tuning paradigms. For fair comparison, all methods are grouped according to the pre-trained trackers. Notably, OneTracker [1] and MamTrack [11] adopt non-standard experimental settings and lack open-source implementations. OneTracker uses a $\times384$ pre-trained OSTrack, incompatible with standard benchmarks, so we compare it separately in Tab. III. MamTrack, with unclear pretraining details but sharing the HiViT-B/16 backbone with SUTrack, is included in SUTrack-based comparisons.
@@ IV-B Comparison with State-of-the-Art Methods
Extensive comparative analyses are presented in Tab. I, Tab. II and Tab. III, where our method demonstrates excellent performance on all multi-modal benchmarks after incorporating the proposed regularized tuning strategies. The corresponding precision and success plots are illustrated in Fig. 6, further substantiating the quantitative results. Evidently, we can observe that both the RGB-only and the cross-modal trackers are becoming increasingly profitable with pre-trained models. In particular, cross-modal approaches exhibit substantial performance gains, highlighting the complementarity between RGB and auxiliary data under complex conditions. Crucially, the remarkable improvements achieved by our method suggest the significance and necessity of developing tailored cross-domain transfer techniques for multi-modal object tracking.
@@
Results on RGB-Event Tracking. As shown in Tab. I and Tab. III, our method surpasses all state-of-the-art trackers across all RGB-Event datasets, achieving the highest precision scores of $98.2\%$, $81.5\%$ and $85.6\%$ on the FE108, VisEvent, and CoeSot datasets, respectively. In particular, on FE108, a dataset characterized by low-light conditions and heavy reliance on event information, our approach surpasses the previous best by a notable margin: $3.0\%$ in PR and $2.4\%$ in SR over OSTrack-B256. The full fine-tuning approaches (e.g., CEUTrack, MamTrack) yield limited improvements, while parameter-efficient fine-tuning paradigms (e.g., ViPT, SDSTrack) face performance bottlenecks. This stems from a misfit that impedes cross-modal transfer, emphasizing the effectiveness of our regularized tuning.
@@
Results on RGB-Depth Tracking. As depicted in Tab. II and Tab. III, our method outperforms all previous state-of-the-art trackers on DepthTrack, obtaining the top performance $67.1\%$ in F-score. Using the pre-trained OSTrack-B256, our method yields substantial improvements: $2.8\%$ in Pr, $4.5\%$ in Re, and $3.7\%$ in F-score. Similarly, when built on the DropTrack-B384 with richer pre-trained knowledge, our method demonstrates superior performance gains, $3.9\%$ in Pr, $4.9\%$ in Re, and $4.4\%$ in F-score. Furthermore, despite no training on the VOT-RGBD2022 dataset, our method still delivers enhanced performance with a $2.3\%$ gain in EAO over existing baselines. This clearly demonstrates the strong generalization and transferability of our approach in cross-modal tracking scenarios.
@@
Results on RGB-Thermal Tracking. As listed in Tab. II and Tab. III, our method achieves $70.1\%$ precision and $56.3\%$ success when evaluated on the pre-trained OSTrack-B256. Furthermore, our method effectively unleashes the potential of the DropTrack, yielding substantial improvements of $4.6\%$ in PR, $3.7\%$ in SR. Remarkably, with SUTrack as our pre-trained baseline, our method sets a new state-of-the-art, reaching $77.8\%$ in PR and $62.9\%$ in SR. Beyond the established benchmarks, our method also excels on the unseen RGBT234 dataset, attaining $93.3\%$ MPR and $70.3\%$ MSR. These results not only underscore the superior performance of our method but also validate its exceptional cross-dataset generalization capabilities.

### v2

@@ III-B Modeling Sensitivity for Multi-Modality Tracking
Manifold hypothesis. As shown in Fig. 3, the parameter set $\{\theta\}_{L=L_{c}}$, with certain loss $L_{c}$, does not fill the entire space $\mathbb{R}^{|\theta|}$. Instead, it usually lies in a manifold $\mathcal{M}_{L_{c}}$, which is also continuous in $\mathbb{R}^{|\theta|}$ [13, 40].
Based on the aforementioned manifold hypothesis, to avoid the loss of generalization ability for fine-tuned models on downstream domains, the critical thing lies in optimizing the model in local tangent space, shown as the red arrow of Fig. 3. Thus, in the following section, we investigate the method to approximate and optimize the model on such tangent space.
@@ III-B1 Safeguarding Generalization via Prior Sensitivity
This generalization gap is measured as the weight distance within the Riemannian manifold defined by FIM. Through geometric perspective, FIM yields insights that large deviations from pre-trained tangent space entail a higher risk of generalization degradation [30, 50]. Therefore, FIM serves as a natural and principled prior sensitivity to reflect the erosion of pre-trained knowledge.
Figure 4: Operation-wise prior sensitivity (i.e., eigen-decomposition FIM map) of the pre-trained OSTrack. The high prior sensitivity indicates a tendency to deviate from the pre-trained tangent space, reflecting the disruption of pretrained knowledge.
Eigen-decomposition Approximation of FIM. Directly computing FIM for a large model is computationally intractable due to its $\mathcal{O}\left(|\theta|^{2}\right)$ complexity. Instead, we propose to approximate its eigen-decompositionwithout explicitly forming the full matrix. Intuitively, this approximation can be likened to analyzing the terrain of the pre-trained loss landscape. Eigenvectors associated with large eigenvalues represent “steep cliffs” directions where modifying parameters would drastically increase the pre-trained loss and destroy prior knowledge. Conversely, small eigenvalues represent “flat valleys” regions where parameters can be adjusted freely to adapt to new modalities without harming the foundation model. By focusing on leading eigenvalues, we efficiently identify the most critical parameter subspaces that must be preserved, while reducing overall complexity.
Specifically, we partition the model’s parameters into $N$ disjoint operation groups, $\theta=\left\{\theta^{1},\cdots,\theta^{N}\right\}$ (e.g. MLPs of FFNs and attention QKVs). Assuming inter-operation independence, we yield a group-diagonal FIM: $\mathcal{F}^{(\theta_{0})}=\mathrm{diag}(\mathcal{F}^{(\theta_{0}^{1})},\dots,\mathcal{F}^{(\theta_{0}^{N})})$, where $\mathcal{F}^{(\theta_{0}^{j})}\in\mathbb{R}^{|\theta^{j}|\times|\theta^{j}|}$ corresponds an independent parameter group $j$. Parameters from different groups thus span orthogonal subspaces with no second-order coupling.
Empirical studies [14, 37] have shown that the FIM of deep networks is spectrally concentrated, featuring a few large eigenvalues and a long, flat tail, which motivates a low-rank approximation using its leading eigenpairs. Let $\tilde{\mathcal{F}}^{(\theta_{0}^{j})}={(V^{j})}\Lambda^{j}{(V^{j}})^{T}$ be the eigen-decomposition, with unit eigenvectors $V^{j}=[v^{j}_{1},\dots,v^{j}_{K}]$ and eigenvalues $\Lambda^{j}=\mathrm{diag}(\lambda^{j}_{1},\dots,\lambda^{j}_{K})$ sorted $\lambda^{j}_{1}\geq\cdots\geq\lambda^{j}_{K}\geq 0$.
@@
See Appendix I-A.
Prior Sensitivity Measure. Specifically, we estimate large eigenvalues of FIM using symmetric finite-difference probes [23]. Around $\theta_{0}$ on $D_{0}$, small fixed-radius perturbations expose the directions that maximize the normalized generalization gap, which is equivalent to the Rayleigh quotient of FIM. This procedure systematically maps the pre-trained loss landscape to recover the sharpest generalization shifts and identify the most sensitive parameters. For each group $\theta^{j}$, by sampling various unit directions $\epsilon^{j}\in\mathbb{R}^{|\theta^{j}|}$ from an isotropic Gaussian distribution $\epsilon^{j}\sim\mathcal{N}\left(\mathbf{0},I\right)$ on the weight sphere, we obtain the empirical measures:
@@ III-B2 Stabilizing Adaptation via Transfer Sensitivity
Figure 5: Instantaneous transfer sensitivity on the VisEvent dataset observed during fine-tuning. The high sparsity in the sensitivity map indicates that only a few gradients (light areas) dominate the updates.
@@
where $\delta$ is a random perturbation centered on the current update step, i.e., $\delta\sim\mathcal{N}\left(\alpha\mathcal{G},\alpha^{2}\sigma^{2}I\right)$, $\alpha$ is the learning rate and $\sigma$ controls the noise. This denotes that updates follow the gradient with inherent noise (due to stochastic sampling, batch gap, etc.).
@@
$\varepsilon_{ada}\leq\alpha\sqrt{2\sigma^{2}}\left\|\mathcal{G}\right\|_{2}.$ (12)
Derivations details of Eq. (9)-(12) can be found in Appendix I-B. Thus, we can deduce that sparse gradients exacerbate tuning instability and adaptation risk. This highlights the need to avoid concentrating updates on specific parameters and emphasizes the importance of a balanced transfer gradient.
@@
$s_{n}^{t}=\left(\frac{\partial\mathcal{L}\left(\theta\mid M\right)}{\partial\theta_{n}}\right)^{2}.$ (13)
@@
Input: pre-trained dataset $D_{0}$, downstream task dataset $D_{t}$, pre-trained weight $\theta_{0}$; number of eigenvalues $K$, number of model operations NN, training steps $T$; initialized eigenvalue set $\Lambda$;
Output: Optimal multi-modal tracker parameters $\theta_{t}$;
Step1: Prior Sensitivity Estimation:
@@
 Calculate prior sensitivity $s_{j}^{p}$ via Eq. (8);
Step2: Transfer Sensitivity Estimation and Regularized Tuning:
@@
  Compute loss $\mathcal{L}(\theta\mid M_{i})$ and gradients;
  Update transfer sensitivity by $s_{n}^{t}=(\frac{\partial\mathcal{L}\left(\theta\mid M_{i}\right)}{\partial\theta_{n}})^{2}$;
  Normalize and combine sensitivity $s_{n}$ by Eq. (14);
  Parameter update via Eq. (15);
@@
Algorithm 1 Sensitivity-Regularized Fine-Tuning
III-C Sensitivity-regularized Tuning
We regularize the learning process by previously discussed sensitivity metrics to derive enhanced trackers. To align with the unified regularization tuning, we rank both the operation-wise prior sensitivity and parameter-wise transfer sensitivity, and then normalize them within the continuous range of $[0,1]$. To harmonize these two sensitivities, we devise a dynamic linear schedule to adjust their weighted combination. At the beginning of training, prior sensitivity plays a dominant role, contributing a weight of $\kappa$ (where $\kappa\in[0,1]$). As training progresses, the influence of transfer sensitivity gradually increases, reaching the same weight $\kappa$ by the end. This ensures that the model initially focuses on retaining pre-trained knowledge and progressively shifts to emphasize training stability. Formally, at each training step $i$, the combined sensitivity is updated as follows:
$s_{n}=\left(\kappa+(1-2\kappa)\frac{i}{T}\right)s_{j}^{p}+\left(1-\kappa-(1-2\kappa)\frac{i}{T}\right)s_{n}^{t},$ (14)
@@ III-C Sensitivity-regularized Tuning
where $\alpha$ is the learning rate. This formulation suggests that more sensitive parameters should retain their previous states to a greater extent, to avoid oscillations or over-adjustments.
The detailed regularization tuning process is given in Algorithm 1. We set the sensitivity harmony coefficient $\kappa=0.6$ in our experiments. Further comprehensive analysis can be found in Table VI.
@@ III-D Learning Objectives
Figure 6: Precision and success plots of the FE108, VisEvent, CoeSot and LasHeR datasets. Zoom in to see details.
@@ IV-A2 Evaluation Metrics
Adhering to recognized standards [17, 51, 16, 41], we evaluated tracking performance with the following metrics. For all RGB-Event benchmarks (FE108, VisEvent and CoeSot), We utilize two widely used metrics, i.e., success rate (SR), precision rate (PR). For DepthTrack, precision (Pr) and recall (Re) are used, with F-score ($F=\frac{2Re\cdot Pr}{Re+Pr}$) as the primary measure. For VOT-RGBD2022, we assess accuracy (Acc), robustness (Rob), and expected average overlap (EAO). For LasHeR, success rate (SR) and precision rate (PR) are employed. For RGBT234, maximum success rate (MSR) and maximum precision rate (MPR) served as evaluation metrics. Notably, our SR/PR precisely aligns with SR/PR in ViPT and OneTrack, Suc/Pre in SDSTrack, Success/Precision in UnTrack, or AUC/P in SUTrack. For all metrics, the larger, the better.
@@ IV-A3 Pre-trained Models and Compared Methods
In this paper, we choose three prototypical one-stream RGB-based trackers as pre-trained baselines: OSTrack [58], the most widely adopted model; DropTrack [49] and SUTrack [9], which offer improved generalization. Notably, OSTrack and DropTrack are built upon ViT-B/16 [12], while SUTrack adopts the HiViT-B/16 [64] as the backbone. Following their pre-trained configurations, these models differ in input resolution: OSTrack ($\times128$$\times256$), DropTrack, and SUTrack ($\times192$$\times384$) for template and search regions, respectively. To comprehensively validate the effectiveness of our method, we conduct the following experiments in Table I and Table II. First, we construct strong RGB-based single-modal baselines under full fine-tuning. Then, we evaluate our method under both full and parameter-efficient fine-tuning paradigms. For fair comparison, all methods are grouped according to the pre-trained trackers. Notably, OneTracker [16] and MamTrack [41] adopt non-standard experimental settings and lack open-source implementations. OneTracker uses a $\times384$ pre-trained OSTrack, incompatible with standard benchmarks, so we compare it separately in Table III. MamTrack, with unclear pretraining details but sharing the HiViT-B/16 backbone with SUTrack, is included in SUTrack-based comparisons.
@@ IV-B Comparison with State-of-the-Art Methods
Extensive comparative analyses are presented in Table I, Table II and Table III, where our method demonstrates excellent performance on all multi-modal benchmarks after incorporating the proposed regularized tuning strategies. The corresponding precision and success plots are illustrated in Fig. 6, further substantiating the quantitative results. Evidently, we can observe that both the RGB-only and the cross-modal trackers are becoming increasingly profitable with pre-trained models. In particular, cross-modal approaches exhibit substantial performance gains, highlighting the complementarity between RGB and auxiliary data under complex conditions. Crucially, the remarkable improvements achieved by our method suggest the significance and necessity of developing tailored cross-domain transfer techniques for multi-modal object tracking.
@@
Results on RGB-Event Tracking. As shown in Table I and Table III, our method surpasses all state-of-the-art trackers across all RGB-Event datasets, achieving the highest precision scores of $98.2\%$, $81.5\%$ and $85.6\%$ on the FE108, VisEvent, and CoeSot datasets, respectively. In particular, on FE108, a dataset characterized by low-light conditions and heavy reliance on event information, our approach surpasses the previous best by a notable margin: $3.0\%$ in PR and $2.4\%$ in SR over OSTrack-B256. The full fine-tuning approaches (e.g., CEUTrack, MamTrack) yield limited improvements, while parameter-efficient fine-tuning paradigms (e.g., ViPT, SDSTrack) face performance bottlenecks. This stems from a misfit that impedes cross-modal transfer, emphasizing the effectiveness of our regularized tuning.
@@
Results on RGB-Depth Tracking. As depicted in Table II and Table III, our method outperforms all previous state-of-the-art trackers on DepthTrack, obtaining the top performance $67.1\%$ in F-score. Using the pre-trained OSTrack-B256, our method yields substantial improvements: $2.8\%$ in Pr, $4.5\%$ in Re, and $3.7\%$ in F-score. Similarly, when built on the DropTrack-B384 with richer pre-trained knowledge, our method demonstrates superior performance gains, $3.9\%$ in Pr, $4.9\%$ in Re, and $4.4\%$ in F-score. Furthermore, despite no training on the VOT-RGBD2022 dataset, our method still delivers enhanced performance with a $2.3\%$ gain in EAO over existing baselines. This clearly demonstrates the strong generalization and transferability of our approach in cross-modal tracking scenarios.
@@
Results on RGB-Thermal Tracking. As listed in Table II and Table III, our method achieves $70.1\%$ precision and $56.3\%$ success when evaluated on the pre-trained OSTrack-B256. Furthermore, our method effectively unleashes the potential of the DropTrack, yielding substantial improvements of $4.6\%$ in PR, $3.7\%$ in SR. Remarkably, with SUTrack as our pre-trained baseline, our method sets a new state-of-the-art, reaching $77.8\%$ in PR and $62.9\%$ in SR. Beyond the established benchmarks, our method also excels on the unseen RGBT234 dataset, attaining $93.3\%$ MPR and $70.3\%$ MSR. These results not only underscore the superior performance of our method but also validate its exceptional cross-dataset generalization capabilities.